Question Part
Points
Submissions Used
A(n) 1.2-kg copper block is given an initial speed of 4.5 m/s on a rough horizontal surface. Because of friction, the block finally comes to rest.
(a) If the block absorbs 85% of its initial kinetic energy as internal energy, calculate its increase in temperature.
°C

(b) What happens to the remaining energy?
It vanishes from the universe.It is so minute that it doesn't factor into the equation.     It becomes chemical energy.It is absorbed by the horizontal surface on which the block slides.
SHOW WORK
Answer #1

mass of copper m = 1.2 kg

Initial speed v = 4.5 m/s

Initial kinetic energy K = (1/2) mv 2

                                 = 12.15 J

Internal energy of the block E = 85 % of K

                                           = 0.85 K

                                           = 10.3275 J

We know E = mC dt

from this increase in temprature dt = E / mC

Where C = 385 J / kg oC

Substitute values you get dt = 10.3275 / (1.2 x 385)

                                          = 0.022 oC

(b).It is absorbed by the horizontal surface on which the block slides